A uniform plane wave of 10 kHz travelling in free space strikes a large block of a
material having ε = 9ε0, μ = 4μ0 and σ = 0 normal to the surface. If the incident
magnetic field vector is given by
B =10^− 3sin (ωt −βy) kˆ
tesla
Write the complete expressions for the incident, reflected, and transmitted field vectors.
1
Expert's answer
2019-09-02T09:20:46-0400

We have frequency "f=10\cdot10^3\text{ Hz}." First, calculate required parameters for each medium (medium of free space has index 0, medium within the block of some material has index 2). Reflected wave has index "r", index "t" is for transmitted wave.


"\beta_0=2\pi f\sqrt{\mu_0\epsilon_0}=\frac{2\pi f}{c}=\frac{\pi}{15000}\text{ m}^{-1},\\
\beta_{r}=-\beta_0,\\
\beta_t=2\pi f\sqrt{\mu_2\epsilon_2}=\frac{\pi}{2500}\text{ m}^{-1},\\
\eta_0=\sqrt{\frac{\mu_0}{\epsilon_0}}=120\pi\space\Omega,\\
\eta_2=\sqrt{\frac{4\mu_0}{9\epsilon_0}}=80\pi\space\Omega,\\
r=\frac{\eta_2-\eta_0}{\eta_2+\eta_0}=-0.2,\\
t=1+r=0.8."


Magnitudes of the incident E-field and H-field:


"E_0=\frac{B_0}{\sqrt{\mu_0\epsilon_0}}=3\cdot10^5\text{ V/m}."

The equations for incident wave:


"\textbf{E}_i=E_oe^{-j\beta_0 z}\hat\textbf{x},\\
\space\\
\textbf{H}_i=\frac{E_o}{\eta_0}e^{-j\beta_0 z}\hat\textbf{y}."


"\textbf{E}_i
=3\cdot10^5\cdot\text{sin}\Big(2\pi\cdot10^4 t-\frac{\pi}{15000}y\Big)(-\hat\textbf{x})\space\frac{\text{V}}{\text{m}},\\
\space\\
\textbf{H}_i=\frac{\textbf{B}_i}{\mu_0}=\frac{2500}{\pi}\cdot\text{sin}\Big(2\pi\cdot10^4 t-\frac{\pi}{15000}y\Big)\hat\textbf{z}\space\frac{\text{V}}{\text{m}}."

The equations for reflected wave with reflection coefficient "r":

"\textbf{E}_r=rE_oe^{-j\beta_r z}\hat\textbf{x},\\
\space\\
\textbf{H}_r=-r \frac{E_o}{\eta_0}e^{-j\beta_r z}\hat\textbf{y}."

"\textbf{E}_r=-6\cdot10^4\cdot\text{sin}\Big(2\pi\cdot10^4 t+\frac{\pi}{15000}y\Big)\hat\textbf{x}\space\frac{\text{V}}{\text{m}},\\
\space\\
\textbf{H}_r=\frac{500}{\pi}\cdot\text{sin}\Big(2\pi\cdot10^4 t+\frac{\pi}{15000}y\Big)\hat\textbf{z}\space\frac{\text{A}}{\text{m}}."

The equations of transmitted wave with transmission coefficient "t":

"\textbf{E}_t=t E_oe^{-\beta_2 z}\hat\textbf{x},\\
\space\\
\textbf{H}_t=t\frac{E_o}{\eta_2}e^{-j\beta_2 z}\hat\textbf{y}."

"\textbf{E}_t=24\cdot10^4\cdot\text{sin}\Big(2\pi\cdot10^4 t-\frac{\pi}{2500}y\Big)(-\hat\textbf{x})\space\frac{\text{V}}{\text{m}},\\
\space\\
\textbf{H}_\tau=\frac{2000}{\pi}\cdot\text{sin}\Big(2\pi\cdot10^4 t-\frac{\pi}{2500}y\Big)\hat\textbf{z}\space\frac{\text{A}}{\text{m}}."